if the earth's population today is 682,386,000,000 and the growth rate of that population is 2% per year, how long will it take(in years) for the population to increase to 10 times its present value
Answer by Alan3354(69443)

if the earth's population today is 682,386,000,000 and the growth rate of that population is 2% per year, how long will it take(in years) for the population to increase to 10 times its present value
---------------------
10 = (1 +0.02)^t (t in years)
1.02^t = 10
t*log(1.02) = 1
t = 1/log(1.02)
t =~ 116.28 years (compounded annually)
-------------------
Population is a continuous process, not just annual adjustments.
10 = e^rt = e^0.02t
0.02t = ln(10)
t = ln(10)/0.02
t =~ 115.13 years, slightly less
115.13 years is the correct answer